OpenStudy (anonymous):

Write the equation of the parabola in vertex form vertex (3,3) point (1,-9) This needs to be f(x)= ???

OpenStudy (anonymous):

\[f(x) = a(x-h)^{2} +k\]

OpenStudy (anonymous):

vertex = (h,k)

OpenStudy (jdoe0001):

so... if you plug in the vertex coordinates given in the template above how would it look like?

OpenStudy (haleyelizabeth2017):

can you help me with a discussion when you are finihed? its the same stuff......:)

OpenStudy (anonymous):

sure

OpenStudy (haleyelizabeth2017):

thanks :)

OpenStudy (anonymous):

f(x)=a(x-3)^2+3?

OpenStudy (anonymous):

yep, but you have to find "a"

OpenStudy (haleyelizabeth2017):

*hint hint~ plug in points

OpenStudy (anonymous):

f(x)=a(1-3)^2+ -9

OpenStudy (haleyelizabeth2017):

you will need to plug in the why as f(x)

OpenStudy (anonymous):

what? I don't understand how to do that

OpenStudy (haleyelizabeth2017):

so you plugged in x, right? Its the same with the y. replace f(x) with -9. so -9=a(1-3)^2+3

OpenStudy (haleyelizabeth2017):

oops....sorry I had put why instead of y....:(

OpenStudy (haleyelizabeth2017):

solve for a. so subtract 3 from y first then square that and add to 3 and then divide both sides by that.

OpenStudy (anonymous):

-9=a(1-3)^2+3 so i would subtract 3 from outside the parenthesis, which would leave me with -9=a(1-3)^2, right?

OpenStudy (haleyelizabeth2017):

it would be-12=a(1-3)^2 -12=a(-2)^2 -12=a(4) -3=a :) you were close

OpenStudy (haleyelizabeth2017):

I kinda worded it wrong at first

OpenStudy (haleyelizabeth2017):

so do you know what the end equation is?

OpenStudy (anonymous):

so i would put -3(-2)^2+3?

OpenStudy (haleyelizabeth2017):

no. it is f(x)=-3(x-3)^2+3

OpenStudy (anonymous):

why is it (x-3) instead of (1-3) being simplified to -2... if that makes sense what I'm saying

OpenStudy (haleyelizabeth2017):

you don't put the points back in. you just have to revert back to the original equation you can up with: f(x)=a(x-3)^2+3

OpenStudy (haleyelizabeth2017):

but you just add the value of a for a.

OpenStudy (anonymous):

oh. that makes a little bit of sense, yeah. thank you

OpenStudy (haleyelizabeth2017):

:) you're welcome!